OpenStudy (anonymous):

What is the area of the shaded portion within the square? Answer: a. 156 m^2 b. 144 m^2 c. 84 m^2 d. 60 m^2 e. Not enough information to answer

OpenStudy (anonymous):

cant see

OpenStudy (anonymous):

Where is the square... (:

OpenStudy (anonymous):

e

OpenStudy (anonymous):

idont know what happen to it

OpenStudy (anonymous):

Hahaha nice amit

OpenStudy (anonymous):

What is the area of the shaded portion within the square?

OpenStudy (anonymous):

its not allowing it to show

OpenStudy (anonymous):

And you have to save and attach it.

OpenStudy (anonymous):

You have to save and attach it... ^_^

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know how to find the area of a triangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. Can you find for me the area of either one of the two triangles inside the square?

OpenStudy (anonymous):

would i multiply the 12 ans 13 together?

OpenStudy (anonymous):

if yes i got 162

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That is the area of the square.

OpenStudy (anonymous):

so how would i find the area of the triangle?

OpenStudy (anonymous):

Well, you should first use Pythagoras theorem to find the other side of the triangle, which would be \(\sqrt{13^2-12^2}=5\). Now the area of each triangle is \(1/2(12)(5)=30\).

OpenStudy (anonymous):

The area of the shaded portion is equal to the area of the square minus the area of the triangles. That's \(12^2-2(30)=84 m^2\)

OpenStudy (anonymous):

how do u get 5 from 13 2 and 12 2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Use Pythagoras theorem. We have a side of the triangle equal to \(12\) and the hypotenuse is \(12\), so the third side is \(\sqrt{13^2-12^2}=5\).

OpenStudy (anonymous):

I meant the hypotenuse is \(13\).

OpenStudy (anonymous):

I'm still having a hard time understanding

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hello anyone there

OpenStudy (anonymous):

remyred90 you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you get an answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

okay, give me the full question again. I will try to help.

OpenStudy (anonymous):

i understand the last stuff he wrote but i didnt get the last part

OpenStudy (anonymous):

I think that you don't have enough information. Wouldn't you need to have the lenght of the shaded area first?

OpenStudy (anonymous):

look at what the person that was helping me was writing and c if u understand plzz

OpenStudy (anonymous):

I see that the answer is 84. It makes sense.

OpenStudy (anonymous):

yea but how did he get 84

OpenStudy (anonymous):

i got the part where he got 5 but afta that idk

OpenStudy (anonymous):

2 triangles times 30 (the area of each)=60, and 12 times 12 is 144. Subtract the 60 from the 144 and you get 84.

OpenStudy (anonymous):

ok let me try this plz give me time to understand before u leave me lol

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

where did the 30 ccum from?

OpenStudy (anonymous):

well, you use the pythagorem therom. a squared plus b squared equalls c squared.

OpenStudy (anonymous):

i really appreciate ur time

OpenStudy (anonymous):

and the area of a triangle = 1/2 times the base times the height of the triangle.

OpenStudy (anonymous):

No problem. Gotta go bell is ringing.

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

i understand it now thank you so much